### v1

# AnyPPG: An ECG-Guided PPG Foundation Model Trained on Over 100,000 Hours of Recordings for Holistic Health Profiling

ABSTRACT
Background: Photoplethysmography (PPG) offers a noninvasive and accessible modality for health monitoring beyond clinical settings. However, existing studies are limited by the scale and diversity of labeled data, constraining model accuracy, generalizability, and the exploration of broader applications. This study investigates the potential of PPG for holistic health profiling through the integration of foundation model techniques.
Methods: We present AnyPPG, a PPG foundation model pretrained on large-scale, multi-source synchronized PPG-electrocardiography (ECG) data. By aligning PPG and ECG representations within a shared embedding space, AnyPPG learns physiologically meaningful features from unlabeled signals. Its capability was further evaluated across a diverse set of downstream tasks, encompassing both conventional physiological analysis and comprehensive multi-organ disease diagnosis, to explore the full potential of PPG in digital health profiling.
Results: Across eleven physiological analysis tasks spanning six independent datasets, AnyPPG achieved state-of-the-art performance, with average improvements of 12.8% in regression and 9.1% in classification tasks over the next-best model. In multi-organ disease diagnosis, AnyPPG demonstrated broad cross-system diagnostic potential. Among 1,014 ICD-10 three-digit disease categories, 13 achieved an area under the receiver operating characteristic curve (AUC) above 0.8 and 137 exceeded 0.7. Beyond strong performance in cardiovascular diseases such as heart failure, valvular disorders, and hypertension, AnyPPG also showed substantial diagnostic value for non-cardiovascular conditions, exemplified by Parkinson’s disease (AUC = 0.78) and chronic kidney disease (AUC = 0.74).
Conclusions: AnyPPG demonstrates that a PPG foundation model trained through physiological alignment with ECG can produce accurate and robust signal representations. Building on this capability, it underscores the potential of PPG as a modality for comprehensive assessment of systemic and multi-organ health. With continued integration into wearable technologies, this approach offers a promising pathway toward precise, scalable, and accessible health monitoring.
INTRODUCTION
Out-of-clinic health monitoring plays an important role in reducing the societal burden of disease and facilitating early detection and prevention. With the rapid proliferation of wearable technologies, physiological signal-based monitoring has emerged as a scalable and cost-effective solution for continuous assessment of health beyond conventional clinical environments 1, 2, 3. Photoplethysmography (PPG), a noninvasive optical technique that measures dynamic changes in peripheral blood volume, offers a practical means to characterize cardiovascular and systemic physiological states 4, 5, 6. Recent advances in deep learning have further enhanced the utility of PPG as an accessible modality for digital health applications, including heart rate estimation 7, hypertension screening 8, atrial fibrillation detection 9, and cardiovascular risk stratification 10.
Despite substantial progress in PPG analysis, most existing studies remain constrained by limited dataset sizes, scarce labeled data, and a narrow scope of downstream tasks. These limitations restrict both model performance and the broader potential of PPG for holistic health profiling, highlighting the need for a unified framework capable of learning robust and transferable representations from large-scale data, together with a systematic investigation into the full extent of PPG capabilities. Recent advances in foundation model technology provide a promising pathway to address these challenges. By pretraining on large, heterogeneous, and multi-source datasets, often through self-supervised learning, foundation models can capture intrinsic patterns within complex physiological data and adapt effectively to diverse downstream tasks 11, 12. Such models have already demonstrated transformative potential across multiple biomedical modalities, including computational pathology 13, 14, 15, echocardiography 16, polysomnography (PSG) 17, 18, and electrocardiography (ECG) 19, 20, 21, 22. Building on these advances, we aim to extend this paradigm to develop an accurate and generalizable PPG foundation model and leverage it to systematically explore the breadth of information that PPG can provide for digital health applications.
In this study, we introduce AnyPPG, a foundation model developed to explore the full potential of PPG in digital health applications. AnyPPG was pretrained on over 100,000 hours of synchronized PPG and ECG recordings from 58,796 participants across five public datasets, enabling the learning of generalizable physiological representations. Unlike previous approaches that relied solely on unimodal PPG data during pretraining 23, 24, 25, AnyPPG performs cross-modal semantic alignment between PPG and ECG representations, enhancing the physiological fidelity and transferability of learned features. This design is motivated by two key considerations. First, ECG captures cardiac electrical activity that is intrinsically coupled with the hemodynamic variations reflected in PPG, providing complementary electrophysiological information that guides representation learning. Second, mounting evidence indicates that multimodal representation learning yields richer and more transferable embeddings than single-modality training 26, 27, 28.
We systematically evaluated the performance of AnyPPG across a broad spectrum of health-related tasks. Using six public datasets, the model was first assessed on eleven conventional physiological analysis tasks, including heart rate estimation and atrial fibrillation detection. Recognizing that PPG reflects peripheral hemodynamics and circulatory dynamics inherently connected to multiple organ systems, we further investigated its potential for multi-organ disease diagnosis using the Multimodal Clinical Monitoring in the Emergency Department (MC-MED) dataset. Specifically, we analyzed 1,014 three-digit International Classification of Diseases, Tenth Revision (ICD-10) categories spanning chapters I-XV to comprehensively characterize the diagnostic landscape of PPG. Across all physiological analysis tasks, AnyPPG consistently outperformed the state-of-the-art model PaPaGei 23 and achieved strong diagnostic performance (area under the receiver operating characteristic curve [AUC] > 0.70) in 137 diseases. Beyond cardiovascular conditions such as heart failure and atrial fibrillation, AnyPPG demonstrated diagnostic utility across diverse non-cardiovascular domains, including Parkinson’s and Alzheimer’s diseases in the nervous system, osteoporosis and arthritis-related disorders in the musculoskeletal system, age-related cataract and glaucoma in ocular diseases, and female genital prolapse and chronic kidney disease in the genitourinary system. These findings indicate that the physiological information captured by PPG extends far beyond cardiovascular applications, underscoring its potential as a scalable biomarker for assessing health across multiple organ systems. The development of AnyPPG provides a unified framework for advancing the understanding of the physiological relevance of PPG across organ systems and lays the groundwork for future research into its broader applications in comprehensive digital health assessment.
Results
Study population and method overview
The pretraining objective of AnyPPG was to align the representations of synchronized PPG and ECG signals within a shared embedding space, enabling the model to learn physiologically grounded and generalizable representations of PPG waveforms. The pretraining was performed using synchronized PPG-ECG recordings from five publicly available datasets: MC-MED 29, PulseDB 30, the Multi-Ethnic Study of Atherosclerosis (MESA) 31, the Human Sleep Project (HSP) 32, and the Cleveland Family Study (CFS) 33. In total, these datasets comprised 109,909 hours of recordings from 58,796 subjects. Detailed descriptions of the pretraining datasets are provided in Table 1. Within MC-MED, only subjects with synchronized PPG-ECG recordings were used for pretraining, whereas the remaining subjects were reserved for downstream multi-organ disease evaluation to ensure complete subject-level independence.
Model performance on conventional physiological analysis tasks was further assessed using six independent public datasets: PPG-DaLiA 34, the Cuff-Less Blood Pressure Estimation Dataset (UCI-BP) 35, the Brno University of Technology Smartphone PPG Database (BUT PPG) 36, the Gyro-Acc-PPG Dataset 37, the Wearable Stress and Affect Detection (WESAD) dataset 38, and DeepBeat 39. For multi-organ disease evaluation, AnyPPG was fine-tuned using the pretraining subset of MC-MED, and its performance was assessed exclusively on the held-out subjects. Detailed dataset descriptions are provided in Table 1.
Table 1: Summary of datasets used for model pretraining and downstream evaluation. R denotes regression tasks, B denotes binary classification, and M-$k$ denotes multiclass classification with $k$ classes. HR, heart rate; SBP, systolic blood pressure; DBP, diastolic blood pressure.
Dataset Used Modality Task Task Type #Subj. (Segments) Recoding Hours
Pretraining
MC-MED PPG & ECG PPG-ECG alignment 49,916 (28,420,140) 78,945
PulseDB PPG & ECG 4,964 (4,596,304) 12,768
MESA PPG & ECG 2,010 (2,860,924) 7,947
HSP PPG & ECG 1,584 (3,333,705) 9,260
CFS PPG & ECG 322 (355,870) 989
Total 58,796 (39,566,943) 109,909
Evaluation
PPG-DaLiA PPG HR estimation R 15 (12,943) 36
UCI-BP PPG SBP estimation R N/A (261,563) 727
DBP estimation R N/A (261,563) 727
BUT PPG PPG HR estimation R 50 (3,840) 11
SBP estimation R 50 (3,840) 11
DBP estimation R 50 (3,840) 11
Signal quality assessment B 50 (3,840) 11
Gyro-Acc-PPG PPG HR estimation R 24 (2,016) 6
WESAD PPG Stress recognition B 15 (4,419) 12
Affect recognition M-4 15 (4,419) 12
DeepBeat PPG Atrial fibrillation detection B N/A (536,399) 1,490
MC-MED PPG Multi-organ disease diagnosis M-1014 15,759 (359,900) 1,000
Indicates an independent test set that was completely excluded from pretraining.
AnyPPG effectively aligns PPG and ECG representations within a shared physiological embedding space
During pretraining, AnyPPG was trained to align PPG and ECG representations within a shared physiological embedding space. Table 2 summarizes the quantitative results of PPG-to-ECG retrieval performance across the held-out test sets of the five pretraining datasets.
Overall, AnyPPG demonstrated strong and consistent cross-modal alignment. When averaged across datasets, the model achieved sample-weighted Recall@1 (R@1), Recall@5 (R@5), and Recall@10 (R@10) scores of 0.736, 0.906, and 0.935, respectively. The mean average precision at 10 (mAP@10) and mean reciprocal rank (MRR) further reached 0.809 and 0.811, reflecting high retrieval accuracy and ranking consistency. Across individual datasets, the PPG-to-ECG retrieval performance remained stable, indicating that the learned representations were both physiologically meaningful and generalizable. Among the datasets, AnyPPG achieved its highest alignment quality on HSP, with R@1, R@5, and R@10 of 0.875, 0.980, and 0.989, and mAP@10 and MRR both at 0.922. The lowest performance was observed on PulseDB, where the model still achieved R@1, R@5, and R@10 of 0.563, 0.870, and 0.941, with corresponding mAP@10 and MRR of 0.692 and 0.695. These results demonstrate that PPG and ECG share coherent and discriminative physiological features, and that AnyPPG effectively captures and aligns this shared information during pretraining.
Table 2: PPG-to-ECG retrieval performance of AnyPPG across datasets. All metrics were computed on batches of 2,560 paired samples and averaged across batches. R@k, recall at k; mAP@10, mean average precision at 10; MRR, mean reciprocal rank.
Dataset #Samples R@1 R@5 R@10 mAP@10 MRR
MC-MED 2,796,347 0.755 0.903 0.926 0.819 0.821
PulseDB 510,408 0.563 0.870 0.941 0.692 0.695
MESA 277,016 0.684 0.913 0.949 0.781 0.783
HSP 333,059 0.875 0.980 0.989 0.922 0.922
CFS 35,854 0.832 0.950 0.969 0.882 0.883
Aggregate (weighted by samples) 3,952,684 0.736 0.906 0.935 0.809 0.811
Aggregate (macro-average) 0.742 0.923 0.955 0.819 0.821
AnyPPG demonstrates superior performance across downstream tasks
AnyPPG was comprehensively evaluated across eleven downstream tasks spanning six datasets, encompassing a broad range of physiological and health-related applications such as vital sign estimation (including heart rate and blood pressure), affect recognition, atrial fibrillation detection, and signal quality assessment. Compared with the baseline models PaPaGei-S and PaPaGei-P, AnyPPG consistently achieved superior performance across all tasks (Figure 1).
Table 3 summarizes the linear probing results for regression tasks. Overall, AnyPPG substantially enhanced predictive accuracy, achieving an average reduction of 12.8% in mean absolute error (MAE) relative to the next-best model. The largest improvements were observed in heart rate estimation, where the MAE reached 9.28, 8.75, and 11.12 beats per minute (bpm) on the PPG-DaLiA, BUT PPG, and Gyro-Acc-PPG datasets, corresponding to relative gains of 32.7%, 9.1%, and 29%, respectively. The coefficient of determination ($R^{2}$) also improved notably, increasing from 0.33, 0.07, and 0.35 to 0.61, 0.19, and 0.60, indicating stronger predictive capability. For classification tasks (Table 4), AnyPPG achieved AUCs of 0.90, 0.82, 0.90, and 0.88 in stress recognition, affect recognition, atrial fibrillation detection, and signal quality assessment, respectively, demonstrating robust discriminative performance. On average, AnyPPG improved AUC by 9.1%, F1-score by 22.6%, and accuracy by 7.8% compared with the next-best model. Together, these results confirm that AnyPPG delivers robust and generalizable performance across diverse physiological analysis tasks.
Figure 1: Performance comparison of AnyPPG and baseline models across downstream tasks under linear probing. Bar plots summarize model performance on both regression and classification benchmarks, with error bars representing standard deviations. MAE, mean absolute error; AUC, area under the receiver operating characteristic curve.
Table 3: Regression results under linear probing. The best result for each task is highlighted in bold. MAE, mean absolute error; $R^{2}$, coefficient of determination.
Dataset Task Model MAE $\downarrow$ $R^{2}$ $\uparrow$
PPG-DaLiA Heart rate estimation PaPaGei-S 13.78 (0.30) 0.33 (0.03)
PaPaGei-P 13.88 (0.13) 0.32 (0.02)
AnyPPG 9.28 (0.18) 0.61 (0.06)
UCI-BP Systolic blood pressure estimation PaPaGei-S 16.26 (0.04) 0.12 (0.00)
PaPaGei-P 15.18 (0.05) 0.21 (0.00)
AnyPPG 13.93 (0.04) 0.32 (0.00)
Diastolic blood pressure estimation PaPaGei-S 8.43 (0.01) 0.06 (0.00)
PaPaGei-P 8.04 (0.03) 0.13 (0.00)
AnyPPG 7.41 (0.03) 0.23 (0.00)
BUT PPG Heart rate estimation PaPaGei-S 10.05 (0.21) 0.01 (0.01)
PaPaGei-P 9.63 (0.37) 0.07 (0.01)
AnyPPG 8.75 (0.29) 0.19 (0.01)
Systolic blood pressure estimation PaPaGei-S 11.81 (0.16) 0.01 (0.01)
PaPaGei-P 11.59 (0.20) 0.04 (0.01)
AnyPPG 11.52 (0.28) 0.06 (0.06)
Diastolic blood pressure estimation PaPaGei-S 6.47 (0.11) 0.00 (0.01)
PaPaGei-P 6.42 (0.20) 0.00 (0.02)
AnyPPG 6.26 (0.12) 0.04 (0.02)
Gyro-Acc-PPG Heart rate estimation PaPaGei-S 19.91 (0.28) 0.03 (0.02)
PaPaGei-P 15.66 (0.27) 0.35 (0.02)
AnyPPG 11.12 (0.65) 0.60 (0.04)
Table 4: Classification results under linear probing. Higher AUC, F1, and accuracy indicate better performance. The best AUC for each task is highlighted in bold. AUC, area under the receiver operating characteristic curve.
Dataset Task Model AUC $\uparrow$ F1 $\uparrow$ Accuracy $\uparrow$
WESAD Stress recognition PaPaGei-S 0.84 (0.01) 0.62 (0.04) 0.80 (0.02)
PaPaGei-P 0.83 (0.02) 0.63 (0.03) 0.80 (0.01)
AnyPPG 0.90 (0.01) 0.79 (0.01) 0.87 (0.01)
Affect recognition PaPaGei-S 0.72 (0.01) 0.42 (0.01) 0.53 (0.01)
PaPaGei-P 0.75 (0.00) 0.45 (0.02) 0.54 (0.02)
AnyPPG 0.82 (0.00) 0.57 (0.01) 0.63 (0.01)
DeepBeat Atrial fibrillation detection PaPaGei-S 0.70 (0.00) 0.48 (0.00) 0.90 (0.00)
PaPaGei-P 0.81 (0.00) 0.60 (0.00) 0.91 (0.00)
AnyPPG 0.90 (0.01) 0.77 (0.00) 0.94 (0.00)
BUT PPG Signal quality assessment PaPaGei-S 0.77 (0.01) 0.64 (0.03) 0.83 (0.02)
PaPaGei-P 0.81 (0.02) 0.69 (0.02) 0.84 (0.01)
AnyPPG 0.88 (0.01) 0.76 (0.02) 0.86 (0.01)
AnyPPG reveals the potential of PPG for comprehensive multi-organ disease diagnosis
Hemodynamic status is closely linked to overall health. Building on this physiological relationship, we further explored the potential of PPG for comprehensive multi-organ disease diagnosis. Specifically, within the MC-MED dataset, AnyPPG was fine-tuned using data from subjects included in the pretraining subset that contained disease annotations. For evaluation, testing was conducted exclusively on data from subjects who were not involved in pretraining. The analysis encompassed 1,014 three-digit disease codes across Chapters I-XV of the ICD-10 classification. To ensure statistical robustness, disease codes with fewer than 100 positive samples in the test set were excluded, yielding a final set of 719 ICD-10 three-digit codes for analysis.
The diagnostic performance of AnyPPG across both disease- and ICD-chapter levels is shown in Figure 2(b). AnyPPG achieved the highest overall diagnostic capability within the circulatory system, while also demonstrating strong discriminative power across multiple other ICD chapters. Among all evaluated conditions, 13 diseases achieved an AUC greater than 0.8, and 137 exceeded 0.7. After excluding non-specific disease descriptions containing terms such as "other", "elsewhere", "not", or "unspecified", 10 diseases retained an AUC above 0.8, and 82 remained above 0.7. The top 50 diseases ranked by diagnostic AUC, along with their descriptions, are presented in Figure 2(a), with the 50th-ranked disease reaching an AUC of 0.73. Within these top 50 diseases, circulatory system disorders were the most common, comprising 19 conditions that included heart failure, valvular diseases (e.g., acute and subacute endocarditis, rheumatic tricuspid valve disorders), arrhythmias and conduction abnormalities (e.g., atrioventricular and left bundle-branch block, atrial fibrillation), and ischemic heart diseases (e.g., chronic ischemic heart disease). The musculoskeletal and connective tissue system contributed seven diseases, primarily involving osteoporosis and arthritis, while the respiratory system included six diseases such as emphysema and pulmonary edema. Beyond these systems, AnyPPG also demonstrated notable diagnostic performance across several others: (1) for neoplastic diseases, elevated AUCs were observed in six cancers, including monocytic leukemia and carcinoma in situ of skin; (2) for eye and adnexa disorders, senile cataract and glaucoma achieved AUCs of 0.76 and 0.74, respectively; (3) for pregnancy, childbirth, and puerperium conditions, hyperemesis gravidarum and early pregnancy hemorrhage both exceeded an AUC of 0.76; (4) for neurological diseases, Parkinson’s disease and Alzheimer’s disease achieved AUCs of 0.78 and 0.77, respectively; (5) for genitourinary disorders, female genital prolapse and chronic kidney disease yielded AUCs of 0.76 and 0.74; (6) for endocrine and metabolic disorders, diabetes mellitus due to underlying disease and amyloidosis both reached an AUC of 0.75; and (7) for infectious and parasitic diseases, cryptococcosis achieved an AUC of 0.74.
For several clinically significant diseases, the results were as follows: (1) for hypertension, the AUCs were 0.74 for primary and 0.73 for secondary types; (2) for diabetes, the AUCs were 0.63 and 0.73 for type 1 and type 2, respectively; (3) for chronic obstructive pulmonary disease, the AUC was 0.76; and (4) for ectopic pregnancy, the AUC was 0.71. Taken together, these findings demonstrate that PPG signals capture physiologically informative patterns reflective of both circulatory and systemic health, highlighting the potential of PPG-based analysis for comprehensive disease assessment.
(a) Top-50 disease-level results ranked by AUC, excluding non-specific diagnostic codes. Bubble size indicates the number of positive samples, and color denotes the corresponding ICD-10 chapter.
(b) Aggregate diagnostic performance of AnyPPG across ICD-10 categories, showing the distribution of AUCs at the disease level (left) and across ICD-10 chapters (right).
Figure 2: Diagnostic performance of AnyPPG across ICD-10 disease categories.
Methods
Pretraining datasets for AnyPPG
Five publicly available datasets were used for the pretraining of AnyPPG, including MC-MED, PulseDB, MESA, HSP, and CFS. Synchronized PPG and ECG recordings from these datasets were utilized for model pretraining, comprising a total of 109,909 hours of data from 58,796 subjects. For each dataset, the paired PPG-ECG recordings were divided into training, validation, and test sets in a ratio of 8:1:1.
MC-MED
The MC-MED dataset 29 comprises 118,385 adult encounters collected at the Stanford Adult Emergency Department between 2020 and 2022. It contains continuously recorded vital signs and physiological waveforms, including PPG, ECG, and respiratory signals. In addition, the dataset provides detailed patient demographics, medical histories, clinical orders, medication records, laboratory and imaging results, and documented clinical outcomes.
PulseDB
The PulseDB dataset 30 is a curated resource developed to benchmark cuffless blood pressure estimation methods. It contains synchronized 10-second segments of PPG, ECG, and arterial blood pressure waveforms from 5,361 subjects, along with demographic metadata such as age. PulseDB comprises two subsets: the MIMIC-III subset (n=2,423) and the VitalDB subset (n=2,938). The MIMIC-III subset includes recordings from patients admitted to the critical care units at Beth Israel Deaconess Medical Center between 2001 and 2012 40, whereas the VitalDB subset contains recordings from surgical patients undergoing routine or emergency procedures at Seoul National University Hospital, Republic of Korea 41.
MESA
The MESA dataset 31 is a longitudinal cohort initiated between 2000 and 2002 to investigate the prevalence and progression of subclinical cardiovascular disease across diverse ethnic populations. Between 2010 and 2012, 2237 participants were enrolled in a Sleep Exam, which includes overnight PSG recordings comprising synchronized channels such as PPG and ECG.
HSP
The HSP dataset 32 comprises 25,941 PSG recordings collected from approximately 19,492 unique patients at the Massachusetts General Hospital Sleep Laboratory as of April 1, 2023. Each recording includes a standardized set of physiological channels, with a subset containing synchronized PPG and ECG signals.
CFS
The CFS dataset 33 is a longitudinal family-based cohort established to quantify the familial aggregation of sleep apnea. Initiated in 1990, the study enrolled 2,284 individuals (46% African American) from 361 families, with participants assessed up to four times over a 16-year period.
Data Preprocessing
All PPG and ECG recordings were preprocessed using a unified pipeline designed to generate temporally aligned and noise-suppressed inputs for pre-training. First, each continuous recording was divided into non-overlapping segments of 10 s duration, and segments containing more than 25% invalid or motionless signals were excluded. The retained segments were then band-pass filtered to remove baseline drift and high-frequency noise: PPG signals were filtered within the range of 0.5-8 Hz following the method of Elgendi et al. 42, while ECG signals were filtered within 0.5-40 Hz and further denoised with a 50 Hz notch filter to eliminate powerline interference. ECG polarity inversion was automatically detected and corrected to ensure consistent waveform morphology. Signal quality was subsequently evaluated for the ECG signals, with quality indices computed according to the criterion proposed by Zhao et al. 43. To enhance the robustness of AnyPPG to noise, no explicit quality screening was applied to the PPG signals. Afterward, the remaining segments were resampled to a uniform sampling rate of 125 Hz and standardized using z-score normalization along the temporal dimension. The resulting synchronized and normalized PPG-ECG segment pairs served as inputs for contrastive pre-training.
Model architecture and pretraining details
The AnyPPG model is built upon the Net1D architecture 44, a one-dimensional convolutional neural network derived from the ResNet framework 45, 46. It consists of two structurally identical encoders that process synchronized PPG and ECG signals to extract modality-specific feature representations. The complete architecture of the encoder and the detailed configuration of its internal modules are summarized in Table 5 and Table 6. Each encoder contains approximately 5.8 million parameters, and transforms an input signal into a 1024-dimensional embedding. Subsequently, two identical projectors map the PPG and ECG embeddings into a shared embedding space. Each projector is composed of a Linear layer, a GELU activation, and another Linear layer, which sequentially reduce the feature dimension from 1024 to 512 and finally to 256. In this 256-dimensional shared space, the embeddings from both modalities are aligned through a CLIP-style contrastive learning framework 26.
Given synchronized PPG-ECG signal pairs $\{(x_{p,i},x_{e,i})\}_{i=1}^{N}$ sampled from the same temporal window, each signal is first encoded by its modality-specific encoder, yielding $z_{p,i}=f_{p}(x_{p,i})$ and $z_{e,i}=f_{e}(x_{e,i})$ (both $\in\mathbb{R}^{1024}$). The encoded features are then mapped by the corresponding projectors to a shared embedding space, $h_{p,i}=g_{p}(z_{p,i})$ and $h_{e,i}=g_{e}(z_{e,i})$ (both $\in\mathbb{R}^{256}$). For stability, we compute cosine similarities on $\ell_{2}$-normalized embeddings (i.e., $h\leftarrow h/\|h\|_{2}$). The model is trained to align the embeddings of corresponding PPG-ECG pairs while pushing apart non-corresponding pairs within the same batch. This objective is implemented through a symmetric InfoNCE loss:
$\mathcal{L}=-\frac{1}{2N}\sum_{i=1}^{N}\left[\log\frac{\exp(\mathrm{sim}(h_{p,i},h_{e,i})/\tau)}{\sum_{j=1}^{N}\exp(\mathrm{sim}(h_{p,i},h_{e,j})/\tau)}+\log\frac{\exp(\mathrm{sim}(h_{e,i},h_{p,i})/\tau)}{\sum_{j=1}^{N}\exp(\mathrm{sim}(h_{e,i},h_{p,j})/\tau)}\right],$
where $\mathrm{sim}(\cdot,\cdot)$ denotes cosine similarity and $\tau$ is a learnable temperature parameter initialized to $0.07$. This bidirectional objective maximizes cross-modal agreement and encourages the encoder-projector networks to map PPG and ECG signals from synchronized cardiac cycles to nearby locations in the shared embedding space, enabling the PPG branch to implicitly learn temporal, morphological, and rhythmic cardiac features that are consistent with ECG physiology.
Table 5: Architecture and parameter summary of the AnyPPG encoder, shared between the PPG and ECG branches. $C_{in}$ and $C_{out}$ denote the input and output channel dimensions. Output shapes are expressed as $[B,C,L]$, where $B$ is the batch size, $C$ the channel dimension, and $L$ the temporal length.
AnyPPG Encoder (Sequential Architecture) $C_{in}$ $C_{out}$ Output Shape $[B,C,L]$ #Parameters
ConvBlock (Type 3) 1 64 [B, 64, 1250] 384
BasicStage2 64 64 [B, 64, 625] 32,064
BasicStage2 64 160 [B, 160, 313] 156,768
BasicStage2 160 160 [B, 160, 157] 172,320
BasicStage3 160 400 [B, 400, 79] 1,413,720
BasicStage3 400 400 [B, 400, 40] 1,510,200
BasicStage1 400 1024 [B, 1024, 20] 2,565,408
MeanPool1d 1024 1024 [B, 1024]
Total 1 1024 [B, 1024, 20] 5,850,864
Table 6: Architectural definitions of the modules used in the AnyPPG PPG and ECG encoders. $C_{in}$ and $C_{out}$ denote the input and output channel dimensions, respectively.
Module Constituent operations (sequential definition)
BasicStagen ($C_{in},\,C_{out}$) BasicBlock (Type 1) ($C_{in},\,C_{out}$) BasicBlock (Type 2) ($C_{out},\,C_{out}$) $\times(n{-}1)$
BasicBlock (Type 1) ($C_{in},\,C_{out}$) ConvBlock (Type 1) ($C_{in},\,C_{out}$) ConvBlock (Type 2) ($C_{out},\,C_{out}$) ConvBlock (Type 1) ($C_{out},\,C_{out}$) Squeeze and excitation attention (reduction ratio=2) MaxPool1d (kernel = 2, stride = 2)
BasicBlock (Type 2) ($C_{in},\,C_{out}$) ConvBlock (Type 1) ($C_{in},\,C_{out}$) ConvBlock (Type 2) ($C_{out},\,C_{out}$) ConvBlock (Type 1) ($C_{out},\,C_{out}$) Squeeze and excitation attention (reduction ratio=2)
ConvBlock (Type 1) ($C_{in},\,C_{out}$) BatchNorm1d($C_{in}$) Swish() Dropout(p=0.5p{=}0.5) Conv1d($C_{in}$, $C_{out}$, kernel_size=1, stride=1, padding=0)
ConvBlock (Type 2) ($C_{in},\,C_{out}$) BatchNorm1d($C_{in}$) Swish() Dropout(p=0.5p{=}0.5) Conv1d($C_{in}$, $C_{out}$, kernel_size=3, stride=1, padding=1)
ConvBlock (Type 3) ($C_{in},\,C_{out}$) Conv1d($C_{in}$, $C_{out}$, kernel_size=3, stride=1, padding=1) BatchNorm1d($C_{in}$) Swish()
Implementation details
Pre-training was conducted on four NVIDIA H20 GPUs, with each GPU processing a batch size of 1536. The model was optimized using the AdamW optimizer 47, with an initial learning rate of $5\times 10^{-4}$, a weight decay of $1\times 10^{-2}$, and a cosine learning rate schedule. Training was performed for a total of 200,000 steps, including 40,000 warm-up steps. Gradient clipping with a maximum norm of 1.0 was applied to stabilize optimization. Both the PPG and ECG encoders were randomly initialized and jointly optimized from scratch. The final checkpoint corresponding to the best validation contrastive loss was selected.
Model evaluation on downstream tasks
Datasets for downstream evaluation
Downstream evaluation was performed across six publicly available datasets, including PPG-DaLiA 34, UCI-BP 35, BUT PPG 36, the Gyro-Acc-PPG dataset 37, WESAD 38, and DeepBeat 39. These datasets collectively cover eleven dataset-task pairs, encompassing heart rate estimation, systolic and diastolic blood pressure estimation, signal quality assessment, stress recognition (binary), affect recognition (multi-class, four levels), and atrial fibrillation detection. In addition, downstream evaluation for multi-organ disease diagnosis was conducted using the MC-MED dataset. For this task, AnyPPG was fine-tuned using only the subset of MC-MED data that overlapped with the pre-training corpus, while the evaluation strictly relied on subjects and recordings that were entirely unseen during pre-training (including all train, validation, and test partitions) to ensure subject-level independence, with 20 PPG segments randomly sampled from each hospitalization record.
Linear probing and fine-tuning strategy
For conventional physiological analysis tasks, linear probing was conducted to assess the quality and linear separability of the learned representations. For each sample, the embedding produced by the frozen encoder was extracted and evaluated on both classification and regression tasks using a nested five-fold cross-validation protocol. For the classification task, a logistic regression model was employed, with inner five-fold cross-validation used to optimize hyperparameters based on macro AUC (one-vs-rest). The search space was defined as $C\in\{10^{-6},10^{-5},\dots,10^{6}\}$, $\text{solver}\in\{\text{lbfgs},\text{saga}\}$, and $\text{penalty}\in\{\text{l2}\}$. For the regression task, ridge regression was adopted, and the inner cross-validation minimized MAE to determine the optimal hyperparameters, with $\alpha\in\{10^{-6},10^{-5},\dots,10^{6}\}$ and $\text{solver}\in\{\text{auto},\text{cholesky},\text{sparse\_cg}\}$. All results were averaged across the five outer folds to provide a reliable estimate of model generalization under linear evaluation. For multi-organ disease diagnosis on the MC-MED dataset, the full parameters of AnyPPG were fine-tuned end-to-end, initialized from the pretrained checkpoint. Fine-tuning was performed separately for each ICD-10 chapter, with each model trained as a multi-label classification task corresponding to the set of diseases within that chapter.
Discussion
PPG offers a noninvasive and highly scalable means of monitoring human physiological health, particularly when integrated with wearable devices. However, its full potential remains underexplored. Existing models are limited by constrained dataset sizes and a lack of diversity in data sources, resulting in suboptimal accuracy and generalizability. Furthermore, most previous studies have focused primarily on traditional physiological monitoring tasks, with insufficient attention to the hemodynamic information captured by PPG signals and their relevance to systemic, multi-organ health. There is a need to comprehensively investigate the potential of PPG through the development of accurate and generalizable models.
In this study, we propose AnyPPG, a PPG foundation model pretrained on large-scale, multi-source synchronized PPG-ECG data. By aligning PPG and ECG representations within a shared physiological embedding space, AnyPPG learns physiologically meaningful and transferable representations of PPG signals. Unlike previous foundation models such as PaPaGei 23, GPT-PPG 24, and PulsePPG 25, which relied solely on unimodal PPG data during pretraining, AnyPPG leverages cross-modal physiological alignment to enhance its capacity for modeling cardiovascular dynamics and improving representation generalizability. Building on this foundation, we systematically evaluated AnyPPG across a range of downstream tasks, encompassing both conventional physiological analyses (e.g., heart rate and blood pressure estimation) and broader explorations of multi-organ disease diagnosis, to comprehensively uncover the potential of PPG in health monitoring and assessment.
In conventional physiological signal analysis tasks, AnyPPG demonstrated state-of-the-art performance. Across 11 downstream tasks spanning six independent datasets, AnyPPG consistently achieved the best results (Figure 1). In regression tasks (Table 3), AnyPPG reduced the MAE by an average of 12.8% compared with the next-best model, while in classification tasks (Table 4), it improved AUC, F1-score, and accuracy by 9.1%, 22.6%, and 7.8%, respectively, highlighting its strong and generalizable performance across diverse applications. Notably, AnyPPG achieved particularly substantial improvements in heart rate estimation and atrial fibrillation detection. For heart rate estimation, the MAE decreased by an average of 23.6%, and the $R^{2}$ improved by 109% on average. For atrial fibrillation detection, the AUC increased from 0.81 to 0.90, and the F1-score rose from 0.60 to 0.77 compared with the next-best model. These notable gains are likely attributable to the cross-modal alignment between PPG and ECG representations during pretraining, as ECG serves as a gold standard for both heart rate estimation and atrial fibrillation detection, providing richer physiological information and superior signal discriminability.
In the multi-organ disease diagnosis tasks, AnyPPG demonstrated strong cross-system diagnostic capability. Across 1,014 three-digit ICD-10 codes from Chapters I-XV, AnyPPG achieved an AUC greater than 0.8 for 13 diseases and greater than 0.7 for 133 diseases. Overall, PPG exhibited the highest diagnostic performance for circulatory system disorders (Figure 2(b)), accurately identifying conditions such as heart failure, valvular diseases, arrhythmias, conduction abnormalities, and hypertension (Figure 2(a)), underscoring the close physiological coupling between PPG signals and cardiovascular health. Beyond the circulatory system, AnyPPG also demonstrated substantial diagnostic potential across multiple other ICD chapters. Representative examples include neurological disorders (e.g., Parkinson’s disease, Alzheimer’s disease), genitourinary diseases (e.g., chronic kidney disease), endocrine and metabolic disorders (e.g., type 2 diabetes), musculoskeletal and connective tissue disorders (e.g., osteoporosis, arthritis), respiratory diseases (e.g., emphysema, pulmonary edema), neoplasms (e.g., monocytic leukemia, carcinoma in situ of the skin), and eye diseases (e.g., senile cataract, glaucoma). All of these conditions achieved an AUC of at least 0.73, indicating that the hemodynamic information captured by PPG signals reflects physiological states across multiple organ systems and supports cross-organ disease recognition. Importantly, several of these diagnostic findings are physiologically interpretable. For example, chronic kidney disease and diabetes frequently co-occur within the cardiovascular-kidney-metabolic (CKM) syndrome framework 48, 49, which is characterized by intertwined cardiovascular, renal, and metabolic dysfunctions. Given that PPG signals reflect peripheral hemodynamic dynamics, they are well suited to capture circulatory abnormalities associated with these CKM-related conditions. Collectively, these findings support the capacity of AnyPPG to reflect systemic health status and underscore the promise of PPG-based modeling for precision health monitoring and early disease detection.
Despite these promising results, several limitations warrant consideration. First, although AnyPPG was pretrained on large-scale, multi-source synchronized PPG-ECG datasets to promote generalizability and demonstrated robust performance across multiple downstream tasks, the pretraining data primarily originated from clinical environments such as emergency departments, intensive care units, and PSG studies. As a result, further evaluation using more diverse real-world wearable data would be valuable to more comprehensively assess its adaptability in everyday monitoring scenarios. Second, while the study systematically examined the diagnostic potential of PPG for multi-organ diseases, the analysis was based solely on the MC-MED dataset and thus lacks external, multi-center validation. Furthermore, given that MC-MED was developed in an emergency care context, certain diagnostic labels may contain inherent variability, which could modestly influence model evaluation for specific conditions. Finally, this work focused primarily on assessing the diagnostic utility of PPG at a multi-organ level, without extending to mechanistic investigations or broader clinical applications. Future studies could explore the potential of PPG for disease risk prediction, longitudinal health monitoring, and population-level stratification to further advance its clinical relevance.
Overall, AnyPPG demonstrates that a foundation model trained through physiological alignment between PPG and ECG can achieve accurate and robust signal understanding. Building on this capability, AnyPPG underscores the promise of PPG as a versatile modality for comprehensive assessment of multi-organ and whole-body health. With further integration into wearable technologies, this approach holds promise for enabling more precise, comprehensive, and accessible personal health monitoring.
Data Availability
All datasets used in this study are publicly available. The MC-MED dataset can be accessed via PhysioNet at https://physionet.org/content/mc-med/1.0.1/. PulseDB is available at https://github.com/pulselabteam/PulseDB. The HSP dataset can be obtained through the Brain Data Science Platform at https://bdsp.io/content/hsp/2.0/. The CFS and MESA datasets are accessible through the National Sleep Research Resource at https://sleepdata.org/. Access to certain datasets (including MC-MED, HSP, CFS, and MESA) requires registration and compliance with the corresponding data usage agreements.
Code Availability
The model implementation and pretrained weights of AnyPPG are publicly available at https://github.com/Ngk03/AnyPPG.
Competing interests
The authors declare no competing interests.
Acknowledgement
This work was supported by the Beijing Natural Science Foundation (QY23040), the National Natural Science Foundation of China (62102008), and the Research Project of Peking University in the State Key Laboratory of Vascular Homeostasis and Remodeling (2025-SKLVHR-YCTS-02).